What is potential Difference?

Grade:12th pass

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear amarnath

The potential difference between any two points A and B in an ΔV =VB -VA
electric field is defined as the change in potential energy of the system divided by
the test charge qo :

    ΔV=ΔU /qo  = -AB E.ds

Potential difference should not be confused with difference in potential energy.
The potential difference is proportional to the change in potential energy, and we
see from Equation that the two are related by
ΔV=ΔU /qo

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin